soving for x
x^2 - 2x >=3

Answer by mananth(16946)

x^2 - 2x >=3
add -3 to both sides
x^2-2x-3>=3-3
x^2-2x-3>=0
x^2-3x+x-3>=0
x(x-3)+1(x-3)>=0
(x-3)(x+1)>=0
either x-3 >=0
x>=3
OR
x+1>=0
x>=-1